Mariel wanted to avail a cellphone plan that offers a

monthly fee of ₱2,500.00. Itincludes 240 minutes of call and charges ₱7.50 for each additional minute of usage.


1
Expert's answer
2021-12-21T04:04:31-0500

According to the passage, let x be the monthly call time

So the monthly cost function will be

"C(x)=2500+(x-240)*7.5\\
C(x)=700+7.5x"